OpenStudy (anonymous):

find three consecutive integers such that the product of the first and the second is equal to the product -6 and the third

OpenStudy (anonymous):

does anybody have an idea

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so do you know what to do? is there a technique

OpenStudy (anonymous):

One approach is to try choosing a number, and try it with that and the next 2 consecutive integers and see if it works. That may sound like guessing...

OpenStudy (anonymous):

is there another way?

OpenStudy (anonymous):

Another approach/technique is to use variables and set up an equation. Notice that consecutive integers look like this: if first integer is "n", then 2nd integer is "n + 1" and third integer is "n + 2"

OpenStudy (anonymous):

@zordoloom do you know?

OpenStudy (anonymous):

So: find three consecutive integers such that the product of the first and the second is equal to the product -6 and the third product of first and second integers: (n) * (n + 1) product of -6 and third integer: (-6) * (n + 2) Those two products are supposedly equal: n(n+1) = (-6)(n + 2) Multiply everything out, simplify, and solve for "n". Then the three consecutive integers are that "n", and the next two integers after.

OpenStudy (anonymous):

so i would just solve it? how do you know that the 2 "n's" are equal. the two integers could be different

OpenStudy (anonymous):

Say, for example, that I know "n" is equal 2. Then what is the next consecutive integer? It's 3, right? But it's also just "n + 1" And the third consecutive integer is 4, but you can also write it as "n + 2" I only used one variable... the other integers are just expressed using that same "n", so the n's are all the same... there's only one "n".

OpenStudy (anonymous):

o ok thanks for clarification. so now i would solve for n(n+1) = (-6)(n + 2)

OpenStudy (aripotta):

how do you solve it? :/ n^2 + n = -6n - 12?

OpenStudy (anonymous):

yes :) Don't worry... I understand your concern about the using the "n" terms. But this is a decent technique that looks better (from a teacher's perspective) than guessing! :)

OpenStudy (anonymous):

what @AriPotta is saying is making sense what should i do next after i distribute

OpenStudy (anonymous):

n(n+1) = (-6)(n + 2) Multiply both sides out first, then collect terms... n(n+1) = (-6)(n + 2) n^2 + n = -6n - 12 n^2 + n + 6n + 12 = -6n + 6n - 12 + 12 (get everything on left) n^2 + 7n + 12 = 0

OpenStudy (anonymous):

then would i n^2+7n=-12 n^2+n =-12/7 ??????????/

OpenStudy (aripotta):

is this...factoring? i haven't learned much about factoring, but this looks like where you would factor :l

OpenStudy (anonymous):

o yeah true i forgot about factoring

OpenStudy (anonymous):

Hmm... no, you need to factor instead.

OpenStudy (anonymous):

i haven't learnt it that well

OpenStudy (anonymous):

ill try

OpenStudy (anonymous):

n^2 + 7n + 12 = 0 What are two numbers you can multiply to get 12 that when you add them, you get 7?

OpenStudy (anonymous):

(n+4) (n+3)

OpenStudy (anonymous):

is that the answer

OpenStudy (aripotta):

no, you still have to find the numbers?

OpenStudy (anonymous):

Yes, (n + 4)(n + 3) = 0 so, that expression is true if n = -4 or if n = -3... so you actually have two values for n.

OpenStudy (aripotta):

:O wat

OpenStudy (anonymous):

The final step should be to check both those solutions for n and see if they work in the original problem. n = -3 next consecutive integer: -2 third consecutive integer: -1 and (-3)(-2) = (-6)(-1) <<--- Yay, true! n = -4 next consecutive integer: -3 third consecutive integer: -2 and (-4)(-3) = (-6)(-2) <<<--- also true, yay! So the integers could be -3, -2, -1 OR -4, -3, -2

OpenStudy (anonymous):

thanks

OpenStudy (aripotta):

ooh

OpenStudy (aripotta):

wait, how'd you know it was negative?

OpenStudy (anonymous):

The factored form was (n + 4)(n + 3) = 0 When you look at those 2 sets of parentheses, ask yourself what n value would make each one = 0? The reasoning is: A * B = 0 if either A or B is = 0, then A * B is equal 0. So, if A = (n + 4) and B = (n + 3) What value of n makes A = 0? What value of n makes B = 0?

OpenStudy (aripotta):

ooooooooooooooh. i get it. thanks :D

OpenStudy (anonymous):

So, @AriPotta, that's why you try to factor in a problem like this... it makes it quick to get the solutions. If you have (x - 5)(x + 6) = 0, then the solutions are x = 5, x = -6... and so on.

OpenStudy (anonymous):

The original equation that would have led to that factoring is: x^2 + x - 30 = 0 (x - 5)(x + 6) = 0 Solutions: x = 5, x = -6

OpenStudy (aripotta):

if i wasn't your fan already, i'd fan ye

OpenStudy (anonymous):

thanks ;)

OpenStudy (anonymous):

A Mathematica solution is attached. The general idea should be apparent.